OpenStudy (anonymous):

help me solve -6k+7k

OpenStudy (anonymous):

-6k + 7k is the same as 7k - 6k Does that help?

OpenStudy (anonymous):

a little but i don't get combining like terms at all

OpenStudy (anonymous):

If you have 7 oranges, and you take 6 oranges away how many oranges do you have?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

1 orange. Yes. So if you have 7 fours. (28) and you take away 6 fours (24) How many fours do you have?

OpenStudy (anonymous):

1 fours i belive

OpenStudy (anonymous):

Indeed. 28 - 24 = 4 = 7(4) - 6(4)

OpenStudy (anonymous):

ok so should wright down this to start solving

OpenStudy (anonymous):

No, just subtract. Think about it this way... You know how to distribute multiplication right? a(b+c) = ab + ac right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what should i be subtracting -6k from 7k

OpenStudy (anonymous):

-6k + 7k = (-6 + 7)k = k